OpenStudy (bohotness):

The city is planning to build a new park. The park will be in the shape of a parallelogram with a perimeter of 568 yards. If the length of the park is 170 yards, what is the width of the park? 114 yards 398 yards 228 yards 3 yards

OpenStudy (loser66):

@Madiyel

OpenStudy (squirrels):

L2+W2=Total amount of yards They already gave us the length so. \[(170\times2)+(W \times 2)=568\]

OpenStudy (squirrels):

L2+W2=Perimeter**

OpenStudy (bohotness):

o.o oh THANK YOU I GOT IT

OpenStudy (squirrels):

No problem :)

OpenStudy (bohotness):

wait

OpenStudy (bohotness):

nvm thank you so much